OpenStudy (anonymous):

The product of two positive numbers is 16. Find the numbers if the sum is least.

OpenStudy (anonymous):

only 3 things can multiply together to equal 16 1 x 16 = 16 4 x 4 = 16 8 x 2 = 16 if those pairs of numbers were added together which would give the smallest sum? 1 + 16 = 17 4 + 4 = 8 8 + 2 = 10 therefore 4 and 4 are the two numbers that give the smallest sum and are both multipliable to 16

OpenStudy (anonymous):

can you show me the equation blackmamba..

OpenStudy (anonymous):

I dont know if there is an equation to solve it in one step but the method I showed you above gives the right answer

OpenStudy (anonymous):

Do you know calculus? You can show that if you optimize a box with side lengths x and y, you get that x=y is the optimized length.